# Constrained Measurement Incompatibility from Generalised Contextuality of Steered Preparation

### v2

@@ Abstract
In a bipartite Bell scenario involving two local measurements per party and two outcome per measurement, the measurement incompatibility in one wing is both necessary and sufficient to reveal the nonlocality. However, such a one-to-one correspondence fails when one of the observers performs more than two measurements. In such a scenario, the measurement incompatibility is necessary but not sufficient to reveal the nonlocality. In this work, within the formalism of general probabilistic theory (GPT), we demonstrate that unlike the nonlocality, the incompatibility of $N$ arbitrary measurements in one wing is both necessary and sufficient for revealing the generalised contextuality for the sub-system in the other wing. Further, we formulate an elegant form of inequality for any GPT that are necessary for $N$-wise compatibility of $N$ arbitrary observables. Moreover, we argue that any theory that violates the proposed inequality possess a degree of incompatibility that can be quantified through the amount of violation. We claim that it is the generalised contextuality that provides a restriction to the allowed degree of measurement incompatibility of any viable theory of nature and thereby super-select the quantum theory. Finally, we discuss the geometrical implications of our results.
@@ I Introduction
It is a distinctive feature of quantum theory over classical theories that the joint probability of measurement outcomes of two observables does not exist, in general. However, if two observables are commuting, they are co-measurable, i.e., the joint probability of them always exits. Therefore, the commutativity is a sufficient condition for the joint measurability of observables. The textbook quantum theory often deals with projective measurements. However, the most general measurement in quantum theory is represented by positive operator-valued measures (POVMs) [1] corresponding to an observable, and the projective value measure is just a special case.
@@
A set of POVMs are called jointly measurable or compatible if there exists a single global POVM that can be suitably marginalized to realize each element of the POVMs from the set. Otherwise, the set of POVMs are called incompatible. Thus, the notion of measurement incompatibility is more general than commutativity, i.e., the latter implies the former but converse does not always hold. Note also that for the case of projective value measure, the pairwise commutativity of a set of arbitrary $N$ number of observables implies the global commutativity of all the $N$ observables. In contrast, the pairwise compatibility of arbitrary $N$ POVMs corresponding to $N$ non-commuting observables does not generally imply the global compatibility of the set [2, 3, 6, 4, 5].
@@
There are many different ways to define the POVMs corresponding to a given observable. However, the unsharp [7, 1] POVMs are commonly discussed ones, which are smeared versions of projective measurements and directly arise from the non-ideal measurement process in quantum theory. Given a set of $N$ unsharp POVMs that are globally incompatible, it can be made compatible if the degree of unsharpness is lower than a critical value [2, 3]. Moreover, the greater the incompatibility of a set of POVMs, the higher the unsharp parameter needs to be taken in order to make the set compatible. Thus, the upper bound to the unsharpness parameter that makes the set of POVMs compatible determines the degree of measurement incompatibility possessed by the concerned measurements.
@@
Over the last two decades, a flurry of important studies have been made to explore the connection between measurement incompatibility and various forms of quantum correlations. It is well-known that measurement incompatibility is necessary to witness Bell nonlocality [9, 10, 8]. Moreover, in the bipartite Bell scenario [10, 8] where each of the observer implements only two dichotomic measurements, it is proved [11, 12] that whenever the measurements performed in one of the wings are incompatible, it is always possible to find shared states and measurements on the other wing such that the joint probability violates the Clauser-Horne-Shimony-Halt inequality. However, for the case of more than two measurements per wing, the measurement incompatibility does not warrant the revelation of nonlocality [13, 14] in general. This qualitatively resembles the fact that for all entangled states, the Bell nonlocality cannot be demonstrated [16, 17, 15]. Thus, measurement incompatibility is necessary but not sufficient to witness Bell nonlocality.
@@
In contrast, it has been recently shown [18] that a form of non-classical quantum correlation - the preparation contextuality [19, 20], can be revealed if and only if the measurements are incompatible. This is closely related to the fact that there exists an one-to-one mapping between measurement incompatibility and a weaker form of nonlocality, the notion of steering [22, 21]. Note that, the non-classical correlations, such as nonlocality and steering, are exhibited for specially separated parties. On the other hand, generalised contextuality [19] refers to a form of nonclassicality that can be revealed through statistics that correspond to a single system.
@@
However, one can leverage generalised contextuality from any of the local systems of a nontrivially correlated bipartite global system by considering that the local measurement on any one of the systems remotely prepares another system in the other spatially separated wing [23]. In particular, for quantum theory, such nontrivially correlated systems are represented as entangled states, which has recently been shown [24] as a prerequisite for obtaining the generalised contextuality of the remotely prepared system. Thus, measurement in one of the wings in a bipartite Bell test can constrict the maximum amount of generalised contextuality in the other wing. In this context, for quantum theory it is already established [22] that in the discussed bipartite scenario the local assemblage of a particular wing is unsteerable if and only if the measurements performed on the other wing are compatible. Supplementing this result, it is also found [18] that if the measurements on one wing are compatible, then the resultant statistics obtained from the systems on other wing must admit preparation non-contextual model and vice-versa.
@@
Considering the growing acceptability of generalised quantum contextuality as a viable notion of nonclassicality behind a plethora of information processing and computational tasks [23, 25, 26, 27, 28, 29, 30, 31], it is important to investigate what structural uniqueness of quantum theory providing such contextual advantage in those tasks. On the other hand, it is also well known that incompatible quantum measurements implies advantage in a number of tasks including quantum state discrimination [32, 33], quantum random access codes [35, 34], and parameter estimation [36]. This indicates that there must be some interrelation between measurement incompatibility and generalised contextuality that are not only significant for quantum theory but also for any other operational theory. Therefore, possible existence of such an interconnection is essential for comprehending the intricacies of quantum theory that singles it out from any other possible no-signalling operational theories.
@@
Against this backdrop, in this work, we investigate this issue from the perspective of generalised probabilistic theory (GPT) [39, 37, 38, 40, 41]. We consider a bipartite Bell scenario involving two parties Alice and Bob, who share a non-trivially correlated system. Alice randomly performs measurements of an arbitrary $N$ number of dichotomic observables in her local subsystem, and Bob randomly performs a certain number of measurements on his part of the system and collects the joint statistics. The primary aim of this paper is to show that in any no-signalling theory, if Alice’s measurements are compatible, then any statistics obtained from Bob’s subsystem admit the generalised non-contextual model. In particular, we argue that, unlike the case of nonlocality, for any no-signalling operational theory, the measurement incompatibility of an arbitrary $N$ number of dichotomic observables in one wing is both necessary and sufficient for revealing the generalised contextuality for the steered subsystems of the other wing. In particular, by invoking the framework of GPT, we demonstrate that the degree of generalised contextuality in a GPT determines the maximum allowed incompatibility that the theory possesses. Furthermore, by noting the geometrical meaning [20] of generalised contextuality we connect the compatibility of measurements at one wing with the geometrical structure of the GPT describing the local statistics of other wing. More precisely, we prove that the local statistics of the steered preparations of Bob admits a simplex-embeddable GPT description if and only if Alice’s measurements are compatible.
@@
Further, in order to establish our key results, we introduce a measure of incompatibility in a GPT by incorporating the fact that the optimal value of fuzziness for which an arbitrary $N$ number of POVMs becomes compatible depends largely on the structure of the no-signalling theory. In this regard, witnesses for pairwise incompatibility of observables in the framework of GPT has already been derived in [42, 43]. Here, we construct an inequality based witness for $N$-wise compatibility valid for any no-signalling probabilistic theory. This witness enables us to quantify the amount of incompatibility of an arbitrary $N$ number of observables corresponding to a particular theory. Having this tools in hand we finally argue that if quantum theory is taken as the governing algorithm of nature then a class of theory whose optimal incompatibility is more or less than quantum theory must not exist. To put it more precisely, by revealing a hitherto unexplored link between two indispensable notions of classicality, namely, the compatibility and the generalised noncontextuality, our result quantifies the maximum degree of measurement incompatibility that nature possesses.
@@
This paper is organized as follows. In Sec. II. we have discussed the operational theories and its abstractions to finally discuss the notion of classicality. In Sec. II we define the incompatibility of measurements in a GPT and discuss the incompatibility of a theory. Sec. IV is equipped with our main arguments and results. In Sec. V we have discussed the optimal incompatibility of measurements allowed in quantum theory while in Sec. VI. we sketched some geometrical implications of our results. e finally conclude with future plans in Sec. VII.
@@ II Ontological model of an operational theory and classicality
Technically speaking, the primitive of an operational theory essentially consists of a set of preparation procedures ${}\{\mathbb{P}\}$, a set of transformations $\{\mathbb{T}\}$, and measurement procedures $\{\mathbb{M}\}$. However, in this paper, we do not explicitly use transformation procedures. Now, given a preparation procedure $P\in\mathbb{P}$ and a measurement procedure $M\in\mathbb{M}$, an operational theory predicts the statistics of the experiment by providing the probability $p(m|P,M)$ of obtaining a particular output $m$. If the operational theory is quantum theory then the preparation procedure is represented by the density matrix $\rho$ and the measurement procedures are in general represented by a set of POVMs $\{E_{m}\}$ where $m$ is the outcome of the measurement. The Born rule provides the probability of obtaining the outcome $m$ of an experiment so that $p(m|P,M)=Tr[\rho E_{m}]$. Although operational theory perfectly predicts the experimental statistics, in general it does not provide an event-by-event causal description behind the statistical predictions. The aim of an ontological model is to provide such a causal description of operational statistics.
@@
It has been shown in the literature that the predictions of quantum theory are in conflict with classical models. This necessitates the articulation of the precise notion of classicality in an ontological model of the operational theory. Certainly, one may impose various assumptions on an ontological model which can lead different forms of classicality. Here, we consider a widely accepted notion of classicality- the generalised non-contextuality [19] .
@@
The notion of generalised non-contextuality can be made precise by introducing equivalence classes of experimental procedures. Let two preparation procedures (measurement procedures) $P_{1}$ and $P_{2}$ ($M_{1}$ and $M_{2}$) cannot be operationally distinguished by any measurement (preparation) for any arbitrary outcome $m$. Hence, the two preparation procedures are operationally equivalent, i.e., $p(m|P_{1},M)=p(m|P_{2},M)$, $\forall M,m$. Similarly, two operationally equivalent measurement procedures satisfy $p(m|P,M_{1})=p(m|P,M_{2})$, $\forall P,m$. In quantum theory, such equivalence is generally represented by preparation procedures that produce the same density matrix $\rho$ and measurement procedures that realize the same POVM element $E_{m}$.
@@
The key question is how such operational equivalences of preparation and measurement procedures in an operational theory can be represented in the ontological model of that operational theory. One may surmise that an equivalent representation of such operational equivalence may also hold in the ontological model. This motivates one to define the notion of generalised non-contextuality [19] which is essentially described in terms of two basic assumptions imposed in any ontological model defined as follows. If any two preparations $P_{1}$ and $P_{2}$ are operationally equivalent then they are represented by the same ontic state probability distributions, i.e., $p(m|P_{1},M)=p(m|P_{2},M)\Rightarrow\mu_{P_{1}}(\lambda)=\mu_{P_{2}}(\lambda)$, $\forall m,M$. This constitutes the assumption of preparation non-contextuality. Along the same line, if two operationally equivalent measurements $M_{1}$ and $M_{2}$ are represented by equivalent response functions in the ontological model, then the model is called measurement non-contextual. Technically, in such a model $\forall m,P,\ p(m|P,M_{1})=p(m|P,M_{2})\Rightarrow\xi_{m|M_{1}}(\lambda)=\xi_{m|M_{2}}(\lambda)$. Simply, a model satisfies both preparation and measurement noncontextual assumptions is called a generalised noncontextual ontological model. The crucial question is up to what extent such a generalised noncontextual ontological model reproduces the statistics of an operational theory. It is already known that certain quantum statistics cannot be reproduced by a generalised noncontextual model., i.e., an ontological model necessarily needs some amount of generalised contextuality in order to compatible with the operational statistics. In this paper, we consider generalised noncontextuality as the notion of classicality to demonstrate our main results in GPT framework.
@@ III GPT and its ontological model
The idea of GPT [37, 38, 39, 41, 40] paves an elegant route towards developing the very framework of possible theories (of the nature) to describe certain observational phenomenon solely from the probabilities. The motivation behind studying the GPTs essentially involves singling out the viable physical theories from a large class of possible theories and pinpointing the underlying reasoning through which one may discard a specific set of among them.
@@
In addition to the description of states, a GPT contains entities that are representative of measurements. For instance, an outcome $m$ of measurement $M$ is represented through the so-called effect $e_{\tiny M}^{\tiny m}\in\Xi$ where $\Xi$ is the effect space. The effect $e_{\tiny M}^{\tiny m}$ is a linear map from a state $\omega$ to a valid probability $p(m|\omega,M)=\langle e_{\tiny M}^{\tiny m},\omega\rangle$ of measurement outcome, i.e., $e_{\tiny M}^{\tiny m}:\Omega\rightarrow[0,1]$ , such that $\forall\omega$ $\sum_{m}\langle e_{\tiny M}^{\tiny m},\omega\rangle=1$. Moreover, there exists a unit measure u for the unique effect u : $\Omega\rightarrow 1$ that corresponds to a measurement with only one outcome that is certain to occur for every state $\omega$. Similar to the embedding of the state space in a vector space $\mathcal{V}$ we will assume that the effect space $\Xi$ will be embedded into the dual of the mentioned vector space defined as $\mathcal{V}^{*}$. Such an embedding is convenient in the sense that one can represent both the state and the effect spaces as hyperplanes and volumes, respectively, embedded into the vector spaces of same dimension. Any valid probability distribution $\langle e_{\tiny M}^{\tiny m},\omega\rangle=p(m|\omega,M)$ is then a bilinear map $\mathcal{M}$: $\mathcal{V}\times\mathcal{V}^{*}\rightarrow\mathbb{R}$.
In general, a GPT describing single system’s states and effects must have the aforementioned properties. However, specific GPTs can be formulated by imposing additional restrictions, which are termed fragments [45] of GPTs. In this paper, we restrict ourselves to this general definition of GPTs and do not assume that every logically possible states and effects are physically realisable, an assumption commonly known as the no-restriction hypothesis [46].

### v3

@@ Abstract
In a bipartite Bell scenario involving two local measurements per party and two outcomes per measurement, the measurement incompatibility in one wing is both necessary and sufficient to reveal the nonlocality. However, such a one-to-one correspondence fails when one of the observers performs more than two measurements. In such a scenario, the measurement incompatibility is necessary but not sufficient to reveal the nonlocality. In this work, within the formalism of general probabilistic theory (GPT), we demonstrate that unlike the nonlocality, the incompatibility of $N$ arbitrary measurements in one wing is both necessary and sufficient for revealing the generalised contextuality for the sub-system in the other wing. Further, we formulate an elegant form of inequality for any GPT that is necessary for $N$-wise compatibility of $N$ arbitrary observables. Moreover, we argue that any theory that violates the proposed inequality possess a degree of incompatibility that can be quantified through the amount of violation. We claim that it is the generalised contextuality that provides a restriction to the allowed degree of measurement incompatibility of any viable theory of nature and thereby super-select the quantum theory. Finally, we discuss the geometrical implications of our results.
@@ I Introduction
It is a distinctive feature of quantum theory over classical theories that the joint probability of measurement outcomes of two observables does not exist, in general. However, if two observables are commuting, they are co-measurable, i.e., the joint probability of them always exits. Therefore, the commutativity is a sufficient condition for the joint measurability of observables. The textbook quantum theory often deals with projective measurements. However, the most general measurement in quantum theory is represented by positive operator-valued measures (POVMs) busch corresponding to an observable, and the projective value measure is just a special case.
@@
A set of POVMs are called jointly measurable or compatible if there exists a single global POVM that can be suitably marginalized to realize each element of the POVMs from the set. Otherwise, the set of POVMs are called incompatible. Thus, the notion of measurement incompatibility is more general than commutativity, i.e., the latter implies the former but converse does not always hold. Note also that for the case of projective value measure, the pairwise commutativity of a set of arbitrary $N$ number of observables implies the global commutativity of all the $N$ observables. In contrast, the pairwise compatibility of arbitrary $N$ POVMs corresponding to $N$ non-commuting observables does not generally imply the global compatibility of the set kunjwal14 ; kunjwal14g ; liang11 ; kunjwal20 ; yu13 .
@@
There are many different ways to define the POVMs corresponding to a given observable. However, the unsharp buschemi24 ; busch POVMs are commonly discussed ones, which are smeared versions of projective measurements and directly arise from the non-ideal measurement process in quantum theory. Given a set of $N$ unsharp POVMs that are globally incompatible, it can be made compatible if the degree of unsharpness is lower than a critical value kunjwal14 ; kunjwal14g . Moreover, the greater the incompatibility of a set of POVMs, the higher the unsharp parameter needs to be taken in order to make the set compatible. Thus, the upper bound to the unsharpness parameter that makes the set of POVMs compatible determines the degree of measurement incompatibility possessed by the concerned measurements.
@@
Over the last two decades, a flurry of important studies have been made to explore the connection between measurement incompatibility and various forms of quantum correlations. It is well-known that measurement incompatibility is necessary to witness Bell nonlocality bell ; chsh ; brunnerrev . Moreover, in the bipartite Bell scenario chsh ; brunnerrev where each of the observer implements only two dichotomic measurements, it is proved wolf09 ; banik13 that whenever the measurements performed in one of the wings are incompatible, it is always possible to find shared states and measurements on the other wing such that the joint probability violates the Clauser-Horne-Shimony-Halt inequality. However, for the case of more than two measurements per wing, the measurement incompatibility does not warrant the revelation of nonlocality bene18 ; hirsch18 in general. This qualitatively resembles the fact that for all entangled states, the Bell nonlocality cannot be demonstrated hororev ; guhnearev ; werner89 . Thus, measurement incompatibility is necessary but not sufficient to witness Bell nonlocality.
@@
In contrast, it has been recently shown tavakoli20 that a form of non-classical quantum correlation - the preparation contextuality spekkens05 ; spekkens21 , can be revealed if and only if the measurements are incompatible. This is closely related to the fact that there exists a one-to-one mapping between measurement incompatibility and a weaker form of nonlocality, the notion of steering uola15 ; quintino14 . Note that, the non-classical correlations, such as nonlocality and steering, are exhibited for specially separated parties. On the other hand, generalised contextuality spekkens05 refers to a form of nonclassicality that can be revealed through statistics that correspond to a single system.
@@
However, one can leverage generalised contextuality from any of the local systems of a nontrivially correlated bipartite global system by considering that the local measurement on any one of the systems remotely prepares another system in the other spatially separated wing schmid18 . In particular, for quantum theory, such nontrivially correlated systems are represented as entangled states, which has recently been shown plavala24 as a prerequisite for obtaining the generalised contextuality of the remotely prepared system. Thus, measurement in one of the wings in a bipartite Bell test can constrict the maximum amount of generalised contextuality in the other wing. In this context, for quantum theory it is already established uola15 that in the discussed bipartite scenario the local assemblage of a particular wing is unsteerable if and only if the measurements performed on the other wing are compatible. Supplementing this result, it is also found tavakoli20 that if the measurements on one wing are compatible, then the resultant statistics obtained from the systems on other wing must admit preparation non-contextual model and vice-versa.
@@
Considering the growing acceptability of generalised quantum contextuality as a viable notion of nonclassicality behind a plethora of information processing and computational tasks schmid18 ; bae21 ; mukherjee22 ; clonning20 ; spekk09 ; pan19 ; saha19 ; schmid22C , it is important to investigate what structural uniqueness of quantum theory providing such contextual advantage in those tasks. On the other hand, it is also well known that incompatible quantum measurements implies advantage in a number of tasks including quantum state discrimination Carmeli19 ; skrzypczyk19 , quantum random access codes carmeli20 ; saha23 , and parameter estimation jae23 . This indicates that there must be some interrelation between measurement incompatibility and generalised contextuality that are not only significant for quantum theory but also for any other operational theory. Therefore, possible existence of such an interconnection is essential for comprehending the intricacies of quantum theory that singles it out from any other possible no-signalling operational theories.
@@
Against this backdrop, in this work, we investigate this issue from the perspective of generalised probabilistic theory (GPT) plavala21 ; janotta11 ; janotta14 ; barrett07 ; hardy01 . We consider a bipartite Bell scenario involving two parties Alice and Bob, who share a non-trivially correlated system. Alice randomly performs measurements of an arbitrary $N$ number of dichotomic observables in her local subsystem, and Bob randomly performs a certain number of measurements on his part of the system and collects the joint statistics. The primary aim of this paper is to show that in any no-signalling theory, if Alice’s measurements are compatible, then any statistics obtained from Bob’s subsystem admit the generalised non-contextual model. In particular, we argue that, unlike the case of nonlocality, for any no-signalling operational theory, the measurement incompatibility of an arbitrary $N$ number of dichotomic observables in one wing is both necessary and sufficient for revealing the generalised contextuality for the steered subsystems of the other wing. In particular, by invoking the framework of GPT, we demonstrate that the degree of generalised contextuality in a GPT determines the maximum allowed incompatibility that the theory possesses. Furthermore, by noting the geometrical meaning spekkens21 of generalised contextuality we connect the compatibility of measurements at one wing with the geometrical structure of the GPT describing the local statistics of other wing. More precisely, we prove that the local statistics of the steered preparations of Bob admits a simplex-embeddable GPT description if and only if Alice’s measurements are compatible.
@@
Further, in order to establish our key results, we introduce a measure of incompatibility in a GPT by incorporating the fact that the optimal value of fuzziness for which an arbitrary $N$ number of POVMs becomes compatible depends largely on the structure of the no-signalling theory. In this regard, witnesses for pairwise incompatibility of observables in the framework of GPT has already been derived in beneduci22 ; Filippov17 . Here, we construct an inequality based witness for $N$-wise compatibility valid for any no-signalling probabilistic theory. This witness enables us to quantify the amount of incompatibility of an arbitrary $N$ number of observables corresponding to a particular theory. Having this tools in hand we finally argue that if quantum theory is taken as the governing algorithm of nature then a class of theory whose optimal incompatibility is more or less than quantum theory must not exist. To put it more precisely, by revealing a hitherto unexplored link between two indispensable notions of classicality, namely, the compatibility and the generalised noncontextuality, our result quantifies the maximum degree of measurement incompatibility that nature possesses.
@@
This paper is organized as follows. In Sec. II. we discuss the operational theories and ontological models to finally discuss the notion of classicality. In Sec. III we describe the GPT and its connection to ontological models. In Sec. IV we define the incompatibility of measurements in a GPT and discuss the incompatibility of a theory. Sec. V contains our main arguments and results. In Sec. VI we discuss the optimal incompatibility of measurements allowed in quantum theory while in Sec. VII. we sketch some geometrical implications of our results and finally conclude with future plans in Sec. VIII.
@@ II Ontological model of an operational theory and classicality
Technically speaking, the primitive of an operational theory essentially consists of a set of preparation procedures ${}\{\mathbb{P}\}$, a set of transformations $\{\mathbb{T}\}$, and measurement procedures $\{\mathbb{M}\}$. However, in this paper, we do not explicitly use transformation procedures. Now, given a preparation procedure $P\in\mathbb{P}$ and a measurement procedure $M\in\mathbb{M}$, an operational theory predicts the statistics of the experiment by providing the probability $p(m|P,M)$ of obtaining a particular output $m$. If the operational theory is quantum theory then the preparation procedure is represented by the density matrix $\rho$ and the measurement procedures are in general represented by a set of POVMs $\{E_{m}\}$ where $m$ is the outcome of the measurement. The Born rule provides the probability of obtaining the outcome $m$ of an experiment so that $p(m|P,M)=Tr[\rho E_{m}]$. Although operational theory perfectly predicts the experimental statistics, in general it does not provide an any causal description of reality behind the statistical predictions. The aim of an ontological model is to provide such a causal description of operational statistics.
@@
It has been shown in the literature that the predictions of quantum theory are in conflict with classical models. This necessitates the articulation of the precise notion of classicality in an ontological model of the operational theory. Certainly, one may impose various assumptions on an ontological model which can lead different forms of classicality. Here, we consider a widely accepted notion of classicality- the generalised non-contextuality spekkens05 .
@@
The notion of generalised non-contextuality can be made precise by introducing equivalence classes of experimental procedures. Let two preparation procedures (measurement procedures) $P_{1}$ and $P_{2}$ ($M_{1}$ and $M_{2}$) cannot be operationally distinguished by any measurement (preparation) for all given outcomes $m$. We call these two preparation procedures to be operationally equivalent, i.e., $p(m|P_{1},M)=p(m|P_{2},M)$, $\forall M,m$. Similarly, two operationally equivalent measurement procedures satisfy $p(m|P,M_{1})=p(m|P,M_{2})$, $\forall P,m$. In quantum theory, such equivalence is generally represented by preparation procedures that produce the same density matrix $\rho$ and measurement procedures that realize the same POVM element $E_{m}$. For explicit examples of states and measurements that are operationally equivalent in quantum theory, we recommend consulting sections IV and V of the ref. spekkens05 .
@@
The key question is how such operational equivalences of preparation and measurement procedures in an operational theory can be represented in the ontological model of that operational theory. One may surmise that an equivalent representation of such operational equivalence may also hold in the ontological model. This motivates one to define the notion of generalised non-contextuality spekkens05 which is essentially described in terms of two basic assumptions imposed in any ontological model defined as follows. If any two preparations $P_{1}$ and $P_{2}$ are operationally equivalent then they are represented by the same ontic state probability distributions, i.e., $p(m|P_{1},M)=p(m|P_{2},M)\Rightarrow\mu_{P_{1}}(\lambda)=\mu_{P_{2}}(\lambda)$, $\forall m,M$. This constitutes the assumption of preparation non-contextuality. Along the same line, if two operationally equivalent measurements $M_{1}$ and $M_{2}$ are represented by equivalent response functions in the ontological model, then the model is called measurement non-contextual. Technically, in such a model $\forall m,P,\ p(m|P,M_{1})=p(m|P,M_{2})\Rightarrow\xi_{m|M_{1}}(\lambda)=\xi_{m|M_{2}}(\lambda)$. Simply, a model satisfies both preparation and measurement noncontextual assumptions is called a generalised noncontextual ontological model. The crucial question is up to what extent such a generalised noncontextual ontological model reproduces the statistics of an operational theory. It is already known that certain quantum statistics cannot be reproduced by a generalised noncontextual model., i.e., an ontological model necessarily needs some amount of generalised contextuality in order to compatible with the operational statistics. In this paper, we consider generalised noncontextuality as the notion of classicality to demonstrate our main results in GPT framework.
@@ III GPT and its ontological model
The idea of GPT janotta11 ; janotta14 ; plavala21 ; hardy01 ; barrett07 paves an elegant route towards developing the very framework of possible theories (of the nature) to describe certain observational phenomenon solely from the probabilities. The motivation behind studying the GPTs essentially involves singling out the viable physical theories from a large class of possible theories and pinpointing the underlying reasoning through which one may discard a specific set of among them.
@@
In addition to the states, a GPT contains entities that are representative of measurements. For instance, an outcome $m$ of measurement $M$ is represented through the so-called effect $e_{\tiny M}^{\tiny m}\in\Xi$ where $\Xi$ is the effect space. The effect $e_{\tiny M}^{\tiny m}$ is a linear map from a state $\omega$ to a valid probability $p(m|\omega,M)=\langle e_{\tiny M}^{\tiny m},\omega\rangle$ of measurement outcome, i.e., $e_{\tiny M}^{\tiny m}:\Omega\rightarrow[0,1]$ , such that $\forall\omega$ $\sum_{m}\langle e_{\tiny M}^{\tiny m},\omega\rangle=1$. Moreover, there exists a unit measure u for the unique effect u : $\Omega\rightarrow 1$ that corresponds to a measurement having only one outcome that is certain to occur given any state $\omega$ from the state space. Similar to the embedding of the state space in a vector space $\mathcal{V}$ we will assume that the effect space $\Xi$ will be embedded into the dual of the mentioned vector space defined as $\mathcal{V}^{*}$. Such an embedding is convenient in the sense that one can represent both the state and the effect spaces as hyperplanes and volumes, respectively, embedded into the vector spaces of same dimension. Any valid probability distribution $\langle e_{\tiny M}^{\tiny m},\omega\rangle=p(m|\omega,M)$ is then a bilinear map $\mathcal{M}$: $\mathcal{V}\times\mathcal{V}^{*}\rightarrow\mathbb{R}$.
In general, a GPT describing single system’s states and effects must have the aforementioned properties. However, specific GPTs can be formulated by imposing additional restrictions, which are termed fragments selby23 of GPTs. In this paper, we restrict ourselves to this general definition of GPTs and do not assume that every logically possible states and effects are physically realisable- an assumption commonly known as the no-restriction hypothesis janotta13 .